OpenStudy (anonymous):

kellie walks 1/4 mile in 1/14 hour. at what rate does kellie walk

OpenStudy (accessdenied):

Do you know this relationship: distance = rate x time ? We know the distance traveled and the time of travel, so we just need to solve for rate.

OpenStudy (accessdenied):

It comes from the idea that, if you travel at a specific rate (say, miles per hour) for an amount of time (hours), the product is the distance traveled (miles/hour * hours = miles). The question gives you a distance and a time: Distance: \( 1/4 \) miles Time: \( 1/14 \) hours So that the relationship is: \( \dfrac{1}{4} = \text{rate} \times \dfrac{1}{14} \)

OpenStudy (accessdenied):

From there, we only need to get rid of the 1/14 multiplied onto the rate to solve for rate. This can be done by dividing both sides by (1/14), or by multiplying both sides by 14.

OpenStudy (anonymous):

mhm

OpenStudy (accessdenied):

In general, this relationship is good to know though: \( \text{distance} = \text{rate} \times \text{time} \) It is always true for average rate, so given any of two of the three variables, you can solve for the last.

OpenStudy (anonymous):

i gott 3 1/2 as my answer

OpenStudy (accessdenied):

I think that looks good. :) rate = 14/4 = 7/2 = 3.5 or 3 1/2

OpenStudy (anonymous):

ok thnks

OpenStudy (accessdenied):

You're welcome! :)